OpenStudy (anonymous):

log3 0.1+2log3x=log3 2+log3 5?

OpenStudy (cwrw238):

2 log3 x = log3 2+log3 5 - log3 0.1 can u use laws of logs to write left side as one log?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i got 10 as an answer. im not sure if that would be right?

OpenStudy (cwrw238):

log3 x^2 = log3 5*2 / 0.1 = log3 100 x^2 = 100 x = 10

OpenStudy (anonymous):

thank you so much :)

OpenStudy (cwrw238):

yw